A computer can do one calculation in 1.4 X 10^-7 seconds. How long would it take the computer to do a trillion (10^12) calculations? (Round to one decimal point)Express answer in scientific notation.
Answer by stanbon(75887)

A computer can do one calculation in 1.4 X 10^-7 seconds. How long would it take the computer to do a trillion (10^12) calculations? (Round to one decimal point)Express answer in scientific notation.
---
Use a proportion
1/1.4x10^-7 = 10^12/x
----
x = 10^12[1.4*10^-7) = 1.4*10^5 seconds
-------------------------------------
=======================================
Cheers,
Stan H.